### v1

# Leveraging Knowledge Graph-Based Human-Like Memory Systems to Solve Partially Observable Markov Decision Processes

Abstract
Humans observe only part of their environment at any moment but can still make complex, long-term decisions thanks to our long-term memory system. To test how an AI can learn and utilize its long-term memory system, we have developed a partially observable Markov decision processes (POMDP) environment, where the agent has to answer questions while navigating a maze. The environment is completely knowledge graph (KG) based, where the hidden states are dynamic KGs. A KG is both human- and machine-readable, making it easy to see what the agents remember and forget. We train and compare agents with different memory systems, to shed light on how human brains work when it comes to managing its own memory systems. By repurposing the given learning objective as learning a memory management policy, we were able to capture the most likely belief state, which is not only interpretable but also reusable. Our code is open-sourced at https://github.com/humemai/agent-room-env-v2-lstm.
1 Introduction
The cognitive science theory suggests that humans have a long-term memory system (Tulving 1983, 1985; Tulving and Thomson 1973). This allows humans to solve various tasks, although they can only observe the world partially. For example, when trying to navigate through a familiar city, humans rely on stored memories of the layout and landmarks. This memory system enables the recall of past experiences, aiding in decision-making and problem-solving even when immediate sensory input is limited. Similarly, when answering a question, humans draw on long-term memory to retrieve relevant facts, concepts, and previous knowledge.
Although there have been many studies, we still lack a comprehensive understanding of how our long-term memory system is managed. Research has identified mechanisms for encoding, storing, and retrieving memories, but the precise regulation and optimization of these processes remain unclear (Artuso and Palladino 2019; Squire 2009).
Understanding memory management is not only a fundamental question in cognitive science but also has significant implications for artificial intelligence (AI). By studying how humans efficiently encode, store, and retrieve information, we can develop more sophisticated AI systems that mimic these processes. AI agents that can manage long-term memory effectively are potentially better equipped to handle complex tasks that require recalling past information, adapting to new situations, and making informed decisions based on incomplete data. This motivation drives our research to create a controlled environment where we can systematically explore these memory management strategies in AI, ultimately aiming to bridge the gap between human cognitive processes and AI (Hassabis et al. 2017; Lake et al. 2017).
The contributions of this paper are as follows.
  • KG-based Environment Design: We design and release a KG-based environment, compatible with Gymnasium (Towers et al. 2023), where an agent has to answer questions while navigating a maze. This environment is highly configurable and offers various levels of complexity.
  • Comparative Analysis of Memory Systems: We compare agents with different memory systems and discuss parallels with how human brains work when it comes to managing its own memory systems.
  • Optimal Memory Management Policy: We show that the best agent is the one that repurposes the given learning objectives as learning a memory management policy, allowing itself to capture the most likely belief state, which is not only interpretable but also reusable.
2 Background
2.1 A Machine With Human-Like Memory Systems
Recent advances in cognitive science and AI have inspired the development of machines with memory systems inspired by human memory. Kim et al. (Kim et al. 2022, 2023) introduced an agent model that explicitly incorporates both semantic and episodic memory systems, demonstrating that such an agent outperforms those with only one type of memory system. Their environment is designed to challenge the agent’s ability to encode, store, and retrieve memories effectively.
Both episodic and semantic memories are part of a human long-term memory system. Episodic memory pertains to the recollection of specific events and experiences that are tied to particular times and places. It involves personal memories that include contextual details about when and where they occurred. Semantic memory, on the other hand, deals with general world knowledge and facts that are not linked to specific occurrences. For example, knowing that Paris is the capital of France is a semantic memory, whereas recalling a visit to the Eiffel Tower involves episodic memory.
To computationally model episodic and semantic memories, Kim et al. (Kim et al. 2022, 2023) add relation qualifier key-value pairs to RDF (Resource Description Framework) triples to make them RDF quadruples (Lanthaler, Wood, and Cyganiak 2014; Hartig et al. 2024). For example, given a memory triple (Agent, atLocation, Library), they would turn this into something like (Agent, atLocation, Library, {timestamp: 42}), and (Agent, atLocation, Library, {strength: 2}), to turn them into episodic and semantic memories, respectively.
KGs are advantageous because they are both human- and machine-readable, facilitating the structured encoding, storing, and retrieving of information.
2.2 Reinforcement Learning (RL) in Partially Observable Markov decision process (POMDP)
RL is useful when supervision of every action is not feasible, such as in memory management. This is often formalized as a Markov Decision Process (MDP) or a Partially Observable MDP (POMDP), which account for partial observability.
We consider an agent interacting with its environment over discrete time steps. The underlying dynamics of the environment is modeled as an MDP (Bellman 1957), but the agent can only partially observe it, making it a POMDP (Åström 1965; François-Lavet et al. 2019). It is defined by: (i) a state space $\mathcal{S}$ that is discrete or continuous; (ii) an action space $\mathcal{A}$ that is discrete or continuous; (iii) the environment’s transition function $T:\mathcal{S}\times\mathcal{A}\to\mathbb{P}(\mathcal{S})$; (iv) the environment’s reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ that is either deterministic or stochastic; (v) an observation space $\Omega$ that is discrete or continuous; (vi) the observation model $O:\mathcal{S}\times\mathcal{A}\to\mathbb{P}(\Omega)$; and (vii) a discount factor $\gamma\in[0,1)$. In our environment, $\mathcal{S}$, $\mathcal{A}$, and $\Omega$, are discrete, $T$ and $O$ are stochastic, and $R$ is deterministic.
The environment starts in an initial state distribution. At time step $t$, the agent maintains a belief state $\mathrm{b}_{t}$, which is a probability distribution over the state space $\mathcal{S}$. The agent then chooses an action $\mathrm{a}_{t}\in\mathcal{A}$ based on its current belief state $\mathrm{b}_{t}$ according to a policy $\pi:\mathcal{B}\rightarrow\mathbb{P}(\mathcal{A})$, where $\mathcal{B}$ denotes the belief space. After taking action $\mathrm{a}_{t}\sim\pi(\mathrm{a}_{t}\mid\mathrm{b}_{t})$, the agent observes an observation $\mathrm{o}_{t+1}\in\Omega$. Using the observation $\mathrm{o}_{t+1}$ and the action $\mathrm{a}_{t}$, the agent updates its belief state to $\mathrm{b}_{t+1}$.
The agent’s objective is to choose a policy $\pi$ that maximizes the cumulative discounted rewards:
$\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$ (1)
where $r_{t}=R(\mathrm{s}_{t},\mathrm{a}_{t})$, $\mathrm{a}_{t}\sim\pi(\mathrm{a}_{t}\mid b_{t})$, $\mathrm{s}_{t+1}\sim T(\mathrm{s}_{t},\mathrm{a}_{t})$, and $\mathrm{o}_{t+1}\sim O(\mathrm{o}_{t+1}\mid\mathrm{s}_{t+1},\mathrm{a}_{t})$. The belief state $b_{t+1}(\mathrm{s}_{t+1})$ is updated using the current belief state $\mathrm{b}_{t}$, the action $\mathrm{a}_{t}$, and the observation $\mathrm{o}_{t+1}$, along with the observation model $O$ and the transition model $T$. The update rule is given by:
$b_{t+1}(\mathrm{s}_{t+1})=\frac{O(\mathrm{o}_{t+1}\mid\mathrm{s}_{t+1},\mathrm{a}_{t})\sum_{\mathrm{s}_{t}\in\mathcal{S}}T(\mathrm{s}_{t+1}\mid\mathrm{s}_{t},\mathrm{a}_{t})b_{t}(\mathrm{s}_{t})}{\sum_{\mathrm{s}^{\prime}\in\mathcal{S}}O(\mathrm{o}_{t+1}\mid\mathrm{s}^{\prime},\mathrm{a}_{t})\sum_{\mathrm{s}_{t}\in\mathcal{S}}T(\mathrm{s}^{\prime}\mid\mathrm{s}_{t},\mathrm{a}_{t})b_{t}(\mathrm{s}_{t})}$ (2)
Updating the belief state can be computationally expensive due to several factors. The belief state represents a probability distribution over the entire state space, which can be vast. Updating it requires summing over all possible states and applying transition and observation probabilities, making the process highly intensive.
A simple way to mitigate this computational expense is to use the history of the past $\tau$ observations $\bm{h}^{(\tau)}_{t}=(o_{t-\tau+1},\ldots,o_{t})$ as the estimated state $\hat{s}_{t}$, with which we learn the policy $\pi(\mathrm{a}_{t}\mid\mathrm{\hat{s}}_{t})$. This alleviates the need to maintain or update the belief state $\mathrm{b}_{t}$.
3 Methodology
3.1 The Rooms Environment
The Rooms Environment provides a simulation for evaluating AI agents with long-term memory systems. Unlike the previous work, this environment consists of multiple interconnected rooms, each containing various types of objects. Static objects, such as beds and tables, remain fixed in their locations and do not move. Independent objects, like Bob and Alice, move between rooms according to probabilistic rules that dictate their movements. Dependent objects, which cannot move by themselves, may be transported by independent objects with a certain probability that depends on the current state of the room. Finally, the agent represents the RL agent that navigates between rooms based on its exploration policy. This configuration allows for the study of how agents manage and utilize long-term memory to make decisions in a dynamic and partially observable environment.
In this environment, hidden states at time $t$, $s_{t}$, are represented as dynamic KGs that capture the entire rooms and the locations of all objects within them. Each room and its contents form a sub-graph of this overall KG. Since the agent can only be present in one room at a time, it can only observe the objects within that particular room. This partial observability requires the agent to utilize its long-term memory.
A dynamic KG is defined as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{T},\mathcal{E})$. Here, $\mathcal{V}$ represents the set of entities, $\mathcal{R}$ the set of relations, $\mathcal{T}$ the set of (discrete) timestamps, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}\times\mathcal{T}$ the set of edges. Each edge in $\mathcal{E}$ is a tuple $(v_{i},r_{k},v_{j},t)$ indicating that a relation $r_{k}$ exists from entity $v_{i}$ to entity $v_{j}$ at time $t$. This structure enables the graph to dynamically reflect the changes within $\mathcal{V}$ and $\mathcal{E}$ over time, under the governance of temporal rules.
Figure 1 shows an example hidden state $s_{t}$ and partial observation $o_{t}$. In this example, there are 32 rooms, i.e., Office room, Kitchen, …, and 25 objects, i.e., Coat, Towel, …, Agent. Therefore, the number of all hidden states is $|S|=N_{\text{rooms}}^{(N_{\text{objects}}-N_{\text{static objects}})}=32^{25-8}\approx 3.8\times 10^{25}$, since every object is located in one of $N_{\text{rooms}}$ and the static objects do not move. The number of possible sub-graphs, i.e., partial observation, that the agent can observe is $2^{(N_{\text{objects}}-N_{\text{static objects}})}\times N_{\text{rooms}}=2^{25-4}\approx 6.7\times 10^{7}$, since the agent can either see or do not see an object in the room. Although the environment might look simple, we have a huge state space, which can be too much for typical table-based RL methods.
In this example, the partial observation (sub-graph of the hidden dynamic KG) is { (Closet, north, Nursery), (Closet, east, Wall), (Closet, south, Wall), (Closet, west, Sauna), (Agent, atLocation, Closet) }. Then the hidden state is a much bigger KG that contains this sub-graph.
(a) An illustration of a hidden state $s_{t}$ (in white) and partial observation $o_{t}$ (in gray).
(b) A hidden state $s_{t}$ (in white) and partial observation $o_{t}$ (in gray) represented as a KG.
Figure 1: An example hidden state $s_{t}$ and partial observation $o_{t}$ of the Rooms environment.
The exploration policy, $\pi^{\text{explore}}(a^{\text{explore}}|o_{t})$, allows the agent to navigate through the room by taking one of five possible actions: North, East, South, West, or Stay. As the agent explores, the environment intermittently poses questions about the object locations. Consequently, the agent must not only learn an effective exploration policy, $\pi^{\text{explore}}$, but also develop a robust question-answering function, $answer=f^{qa}(o_{t})$. The question-answering function also serves as the reward function: the agent receives a reward of +1 for each correctly answered question and 0 otherwise. Although straightforward, using $o_{t}$ for the exploration and question answering is not enough. Therefore, the agent should take advantage of its previous partial observations to form memory.
To maximize its rewards, the exploration policy should guide the agent to spend more time in “informative” rooms, where the likelihood of encountering valuable information is higher. This is crucial because objects do not spend equal amounts of time in each room, meaning that not all sub-graphs are equally informative. Therefore, the agent needs to utilize its long-term memory to maintain and retrieve information about the maze layout and object locations, which is essential for both exploration and answering questions.
In this work, “question answering” is approached as a form of KG reasoning, specifically restricted to answering one-hop queries. These queries involve predicting the tail in a given RDF triple $(s,r,?)$, where $s$ is the subject, $r$ is the relation, and the goal is to predict the unknown tail. This type of reasoning involves traversing the graph from a known entity using a specified relation.
When the environment randomly samples objects (RDF triples) to generate questions, not all objects have an equal probability of being sampled. This bias in sampling affects both the exploration policy and the question-answering function. The agent must account for this sampling bias by focusing its exploration on areas with a higher likelihood of being queried.
The environment randomly samples a triple from its current hidden state and turns it into a question. In our running example, for instance, (Phone, atLocation, Kitchen) and (Alice, atLocation, Library) can be selected and turned into (Phone, atLocation, ?) and (Alice, atLocation, ?), respectively.
The Rooms environment is highly configurable, allowing for a wide range of experimental setups. Researchers can adjust various parameters such as the frequency of questions asked, the total number of questions per episode, the number of rooms, the layout of the rooms, the number and types of objects, and the number of steps per episode, etc. This flexibility enables the creation of diverse scenarios and difficulty levels, facilitating comprehensive studies on the effectiveness of different strategies.
3.2 Agents and Their Learning Objectives
Baseline Agent
The baseline agent utilizes the history of the past $\tau$ observations, denoted as $\bm{h}^{(\tau)}_{t}=(o_{t-\tau+1},\ldots,o_{t})$, to represent the current estimated state $\hat{s}_{t}$. This approach is the most straightforward method for handling a POMDP. For the question-answering function $f^{qa}$, the agent employs a simple strategy: using the most recent relevant memory to answer the question. This strategy assumes that the most recent observation contains the most accurate and relevant information for responding to queries about the environment.
In this setup, solving the POMDP involves optimizing the exploration policy $\pi^{\text{explore}}$, where the reward function the question-answering function $f^{qa}$. Thus, the “current state” for $\pi^{\text{explore}}$ is $\bm{h}^{(\tau)}_{t}$, and the “next state” is the updated $\bm{h}^{(\tau)}_{t+1}$, where the oldest observation is removed, and the new observation is added to the sequence.
HumemAI Agent
The agent with human-like memory systems (hereinafter HumemAI) introduces a function, $f^{\text{mm}}$, known as the memory management function. This agent features a memory system $\bm{M}_{t}$, which includes both short-term and long-term memory components, $\bm{M}^{\text{short}}_{t}$ and $\bm{M}^{\text{long}}_{t}$, respectively. The long-term memory system is further divided into episodic ($\bm{M}^{\text{episodic}}_{t}$) and semantic ($\bm{M}^{\text{semantic}}_{t}$) memory.
The function $f^{\text{mm}}$ first transfers the current partial observation to the short-term memory system by adding the relation qualifier current_time along with its value. For example, (Agent, atLocation, Home Office) at $t=42$ becomes (Agent, atLocation, Home Office, {current_time: 42}). The function then determines whether this short-term memory should be (1) moved to the episodic part of the long-term memory, (2) moved to the semantic part of the long-term memory, or (3) forgotten. If option (1) is selected, the short-term memory is stored in the episodic memory as (Agent, atLocation, Home Office, {timestamp: 42}). If option (2) is chosen, it is stored as (Agent, atLocation, Home Office, {strength: 1}).
The “strength” term indicates the generalizability of this memory in the semantic memory system, with higher values representing repeated additions of the same memory. We also introduce an exponentially decaying factor for each semantic memory to ensure that all memories gradually weaken over time, allowing for continuous adaptation. The decaying of semantic memory is well studied in cognitive science (Rogers and Friedman 2008; Milberg et al. 1999; Silveri et al. 2018; Catricalà et al. 2015).
[…]
We aim to learn the function $f^{\text{mm}}$ using RL. To achieve this, we set the exploration policy $\pi^{\text{explore}}$ using a handcrafted function $f^{\text{explore}}$, which directs the agent to “go to random rooms while avoiding walls, utilizing the long-term memory.” For the question-answering function $f^{qa}$, we use the strategy of “looking up the most recent relevant or the strongest relevant memory.” By doing this, we have repurposed the original learning objective to learning the memory management policy $\pi^{\text{mm}}(a^{\text{mm}}_{t}|\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t})$, where $a^{\text{mm}}_{t}$ represents one of the three memory management actions. As for $\pi^{\text{mm}}$, the “current state” is the current memory and the “next state” is the next memory. Figure 2 shows the four steps involved in training.
Figure 2: A visualization of the four steps involved in training.
The learning process involves two distinct phases. First, we focus on learning $\pi^{\text{mm}}(a^{\text{mm}}_{t}|\bm{M}^{\text{long}}_{t-1},\bm{M}^{\text{short}}_{t})$. This phase ensures that the agent can effectively manage its long-term memory by deciding which observations to store as episodic or semantic memories, or which to forget. Once the memory management function is optimized, we then proceed to the second phase, which involves learning the exploration policy $\pi^{\text{explore}}(a^{\text{explore}}_{t}|\bm{M}^{\text{long}}_{t})$, by taking advantage of the function $f^{\text{mm}}$ learned in the first phase. In this phase, the agent uses its long-term memory to make informed navigation decisions within the environment. Here, the “current state” for $\pi^{\text{explore}}$ is the current long-term memory $\bm{M}^{\text{long}}_{t}$, and the “next state” is $\bm{M}^{\text{long}}_{t+1}$, updated by $\bm{M}^{\text{long}}_{t+1}=f^{\text{mm}}(\bm{M}^{\text{long}}_{t},\bm{M}^{\text{short}}_{t+1},a^{\text{mm}}_{t+1})$
An advantage of this two-phase approach is that the human-like long-term memory can effectively capture the most likely snapshot of the environment (the most likely belief state). This enables the agent to make well-informed decisions on exploration and question-answering.
3.3 The Learning Algorithm
Given the discrete nature of the action spaces for both the baseline and the HumemAI agents, since all of our policies have discrete action spaces, we opt for a value-based RL algorithm (Watkins and Dayan 1992), i.e., DDQN (Hasselt, Guez, and Silver 2016) instead of the vanilla DQN (Mnih et al. 2013), to avoid the overestimation of state-action values. On top of that, we decouple value and advantage in the deep Q-networks (DQN), i.e., dueling network (Wang et al. 2016), to generalize learning across actions. A stochastic gradient descent algorithm, Adam (Kingma and Ba 2015), was used to minimize the temporal difference (TD):
$\mathcal{L}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim D}\left[\left(y-Q(s,a;\theta)\right)^{2}\right]$ (3)
where
$y=r+\gamma Q(s^{\prime},\arg\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta);\theta^{-})$ (4)
In the above equations, $\mathcal{L}(\theta)$ represents the TD loss function. $(s,a,r,s^{\prime})$ are the state, action, reward, and next state, respectively, sampled from the replay buffer $D$. $Q(s,a;\theta)$ is the Q-value predicted by the current Q-network parameterized by $\theta$. $y$ is the target value computed using the reward $r$ and the discounted maximum Q-value of the next state $s^{\prime}$. $\gamma$ is the discount factor. $\theta^{-}$ represents the parameters of the target Q-network, which are periodically updated with $\theta$.
By minimizing this loss, we aim to bring the predicted Q-values closer to the target Q-values, thus improving the policy’s performance over time.
The DQN receives as input either $\bm{M}_{t}$ for HumemAI agent or $\bm{h}^{(\tau)}_{t}$ for the baseline agent, and outputs the Q-values of the actions. This deep neural network involves KG embeddings (KGEs), LSTMs (Hochreiter and Schmidhuber 1997), and MLPs (Sanger and Baljekar 1958). Algorithm 1 shows how a forward pass is made to compute the Q-values for the HumemAI agent.
Algorithm 1 A forward pass to compute the Q-values
Input: A memory system $\bm{M}$; the maximum qualifier value for normalization, $max\_val$; the learnable parameters including the embedding matrix $\bm{E}$, scaling factors for the relation qualifiers $\bm{s}$, and $\bm{\theta}^{\text{LSTM}}$, $\bm{\theta}^{\text{MLP}},\bm{\theta}^{\text{key}},\bm{\theta}^{\text{query}},\bm{\theta}^{\text{value}}$.
Output: The Q-values.
1:  Step 1: Convert Memory System to KGE
[…]
3:     $\operatorname{KGE}(\bm{M}^{(i)})\leftarrow$ [] // initialize as an empty list
4:     $\bm{M}_{\text{sorted}}^{(i)}\leftarrow\bm{M}^{(i)}$ sorted by qualifier value.
5:     for all $(h,r,t,\text{$key$-$value$ pair})\in\bm{M}_{\text{sorted}}^{(i)}$ do
6:        $normalized\leftarrow\frac{value}{max\_val}$
7:        Append $\bm{E}[h]+\bm{E}[r]+\bm{E}[t]+(\bm{E}[key]\times normalized\times\bm{s}[key])$ to $\operatorname{KGE}(\bm{M}^{(i)})$
8:     end for
9:  end for
10:
11:  Step 2: Obtain LSTM-based Representations
[…]
13:     $\bm{h}^{(i)}\leftarrow\operatorname{LSTM}(\operatorname{KGE}(\bm{M}^{(i)}))$
14:  end for
15:  Stack the column vectors $\bm{h}^{(i)}$ horizontally to form the matrix $\bm{H}=[\bm{h}^{\text{short}},\bm{h}^{\text{episodic}},\bm{h}^{\text{semantic}}]^{T}$.
16:
17:  Step 3: Compute Attention Weights
18:  $\bm{Q},\bm{K},\bm{V}\leftarrow$$\hskip 20.00003pt\text{QueryNet}(\bm{H}),\text{KeyNet}(\bm{H}),\text{ValueNet}(\bm{H})$
19:  $\bm{A}\leftarrow\text{softmax}(\bm{Q}\bm{K}^{T})$
20:
21:  Step 4: Attention-weighted Summation
22:  $\bm{v}_{\text{mem}}\leftarrow\left(\sum_{i}(\bm{A}\bm{V})_{i,:}\right)^{T}$
23:
24:  Step 5: Calculate Q-values
25:  $Q(\bm{M},a)\leftarrow\operatorname{MLP}(\bm{v}_{\text{mem}},a;\bm{\theta}^{\text{MLP}})$
26:
27:  return $Q(\bm{M},a)$
The algorithm to compute the Q-values for the HumemAI agent is for one policy only. Since this agent learns two policies, i.e., $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, the algorithm is repeated twice. All the weights are randomly initialized except one: $\bm{\theta}^{\text{LSTM}}_{\text{explore}}$ is initialized with $\bm{\theta}^{\text{LSTM}}_{\text{mm}}$ because we want to leverage the memory management knowledge already learned by the agent.
As for $\pi^{\text{mm}}$, out of $N_{\text{observations}}$ observations within a given room, only one of them gets stored in the short-term memory system at a time. This is because each quadruple (observation) requires an action. In the end, all $N_{\text{observations}}$ of them will be moved to the episodic, semantic, or be forgotten, but they are done sequentially, not all at once.
The algorithm is simpler for the baseline agent, since it does not have different memory types and it just learns one exploration policy. It does not need scaling factors or attention.
KGE converts a symbolic KG into a list of learnable embeddings. Batch processing for the LSTM requires padding to ensure all samples are of equal length. Memories are ordered to help the LSTM discern the relative differences in qualifier values, such as which memories were stored earlier and which have higher strength values. The output of the LSTM is the final hidden representation, which encapsulates the processed information from the sequence of memories. $max\_val$ is the maximum number of steps in an episode, e.g., 100. $\bm{s}$ helps adjust the influence of different qualifiers based on their significance. $\bm{A}$ performs attention to weigh the contributions of different memory types. The dueling network architecture is integrated within the MLP. It consists of two sub-networks: one for computing the value function and another for computing the advantage values. These sub-networks are then combined to calculate the final Q-values. Figure 3 shows the DQN architecture for the HumemAI agent.
Figure 3: Diagram of the DQN architecture for the HumemAI agent, illustrating the steps from memory as KGs to Q-value calculation. The Q-values can be for either memory management actions or exploration actions.
4 Experiments
We evaluate the baselines with varying sizes of $\bm{h}^{(\tau)}_{t}$. Additionally, for the HumemAI agent, we create two extra variants for ablation: one utilizing only an episodic memory system and the other utilizing only a semantic memory system. This allows us to analyze the individual contributions of each memory type to the overall performance.
We designed the training setup to ensure fair comparison between all agents. The HumemAI agent, trained in two phases, has about 85,000 learnable parameters per phase and undergoes 100 training episodes. The decaying factor semantic memory was set to $0.9$. For parity, baseline agents have 170,000 learnable parameters and 200 training episodes. All other hyperparameters remain constant across agents. We select the best-performing model from training and runs in a separate test environment with a greedy policy. We ran in total of 5 runs to report average performance. Table 1 shows the overall results.
Capacity Agent Type Phase 1 Phase 2
12 HumemAI (E) 191 ($\pm$42) 194 ($\pm$29)
HumemAI 105 ($\pm$37) 160 ($\pm$30)
Baseline N/A 144 ($\pm$14)
HumemAI (S) 111 ($\pm$43) 124 ($\pm$65)
24 HumemAI 127 ($\pm$26) 214 ($\pm$64)
HumemAI (E) 227 ($\pm$21) 209 ($\pm$30)
Baseline N/A 138 ($\pm$52)
HumemAI (S) 98 ($\pm$45) 112 ($\pm$79)
48 HumemAI 118 ($\pm$18) 235 ($\pm$37)
HumemAI (S) 192 ($\pm$13) 226 ($\pm$97)
HumemAI (E) 201 ($\pm$42) 225 ($\pm$25)
Baseline N/A 200 ($\pm$15)
96 Baseline N/A 155 ($\pm$77)
192 Baseline N/A 144 ($\pm$68)
Table 1: Average test episode rewards for various agent types and capacities. Standard deviations are written in the parentheses (It was run 5 times with random seeds). HumemAI (E) is the HumemAI agent only with an episodic memory system, while (S) is the one only with semantic. The Phase 1 column represents the phase where we learn $\pi^{\text{mm}}$ while $f^{\text{explore}}$ remains fixed with a heuristic. Phase 2 denotes the phase where we learn $\pi^{\text{explore}}$ while $f^{\text{mm}}$ is the neural network learned from Phase 1. One episode consists of 100 steps, where each step involves 10 random questions, resulting in a total of 1,000 questions. Although a theoretical maximum episode reward is 1,000, this is unattainable in practice, as the agent cannot maintain a complete hidden state in its memory throughout the entire episode.
Table 1 shows the superior performance of the HumemAI agent, especially those with both episodic and semantic memory systems. The two-phase training approach proves effective, with Phase 1 optimizing memory management and Phase 2 leveraging this to improve exploration. The HumemAI agents consistently outperform baselines across memory capacities. Interestingly, longer observation histories for baseline agents don’t necessarily improve performance, as seen with the capacity 192 baseline performing worse than the HumemAI agents with capacity 12. This suggests that longer histories may include irrelevant information. Training efficiency is also notable, with the HumemAI agent (capacity 24) requiring 40 minutes in total, while a baseline agent (capacity 48) took about 70 minutes on a single-CPU machine (AMD Ryzen 9 7950X 16-Core Processor). The exact details of the hyperparameters can be found in the supplementary material.
Figure 4 illustrates the attention weights of $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$ over time for the HumemAI agent with capacity 24. In 4(a), we observe a decrease in the agent’s focus on short-term memory as the agent moves between the rooms (observes different sub-graphs), while attention to long-term memory increases. This shift is logical, given that the long-term memory starts empty at the beginning of an episode. Notably, for both $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, the agent shows a preference for the episodic component of long-term memory. This preference suggests that episodic memories play a crucial role in both memory management and exploration strategies.
(a) $\pi^{\text{mm}}$
(b) $\pi^{\text{explore}}$
Figure 4: Attention weights of $\pi^{\text{mm}}$ and $\pi^{\text{explore}}$, respectively, over time for the HumemAI agent with $capacity=24$
Figure 5 illustrates the agent’s memory systems at the final time step of an episode. It essentially captures information about the hidden belief state. The agent predominantly uses episodic memory for atLocation relations (7 in episodic vs. 4 in semantic), suggesting a preference for timestamped object locations. This strategy likely aids in answering questions by distinguishing between old and new memories. Conversely, map layout information is stored more in semantic memory (8 semantic vs. 5 episodic), reflecting its factual and static nature. Notably, the agent never stores wall-related memories in episodic memory, possibly because the environment never queries wall locations. This memory distribution demonstrates the agent’s ability to adapt its storage strategy based on the nature of the information and its relevance to the task at hand.
Figure 5: An example of the agent’s (the HumemAI agent with $capacity=24$) memory systems, $\bm{M}_{t=99}$. The nodes are colored differently by the memory systems they belong to.
5 Related Work
Our work stands at the intersection of cognitive science, KG, and RL, and while there has been significant research in each of these areas individually, there is relatively little work that combines them to address the challenges of partially observable environments. Notably, the term “episodic memory” is often used in RL to refer to state-action pairs experienced by the agent, similar to the replay buffer used in DQN (Mnih et al. 2013). However, this usage diverges from our approach, which adheres closely to the cognitive science definition of episodic memory introduced by Tulving et al. (Tulving 1983, 1985; Tulving and Thomson 1973), emphasizing its recency property.
The cognitive architectures like ACT-R (Anderson 2007) and Soar (Laird 2012) offer comprehensive frameworks for understanding human memory systems and their applications in AI agents, yet they fall short in providing computational methods for training RL agents with these memory systems. Extensions of these frameworks, such as the work by Gorski and Laird (Gorski and Laird 2011) and Li (Li 2020), focus on learning episodic memory retrieval within these architectures, though they do not leverage KGs or sophisticated RL techniques.
Memory-based RL approaches, such as those by Ming et al. (Meng, Gorbet, and Kulić 2021) and Lin et al. (Lin et al. 2018), add memory components to address partial observability and sample efficiency, respectively. Lin et al. use episodic memory to regularize DQN but differs from our method by not leveraging KGs for interpretability and long-term memory management.
In a similar vein, Han et al. (Han et al. 2019) integrate RL with memory systems to improve question-answering performance. However, their memory system lacks the structured, interpretable nature of KGs, relying instead on numeric embeddings. Demir (Demir 2023) introduces memory-modifying actions and intrinsic motivation to guide memory retention in partially observable environments, focusing on event sequences without employing KGs.
Young et al. (Young, Sutton, and Yang 2018) and Hu et al. (Hu et al. 2021) propose methods for maintaining episodic memory through reservoir sampling and generalizable episodic memory representations, respectively. These approaches aim to improve RL performance by conditioning policy networks on past experiences, but they do not address the structured nature of KGs or the human-like episodic-semantic memory systems utilized in our work.
Incorporating KGs into RL has primarily focused on KG reasoning (Wan et al. 2020; Das et al. 2018; Xiong, Hoang, and Wang 2017; Wang et al. 2022; Park et al. 2022), where the goal is to learn to reason about the relationships between entities. These works typically use current triples or concatenated visited triples as state representations, differing from our approach that constructs and updates a dynamic KG as part of the agent’s state representation. The actions in KG reasoning involve selecting outgoing edges from the current node, whereas our actions include memory management and navigation based on the KG state.
Several works use KGs as prior knowledge to enhance RL performance. Höpner et al. (Hopner, Tiddi, and van Hoof 2022) improve sample efficiency by leveraging KGs, while Piplai et al. (Piplai et al. 2020) guide exploration using KGs. Chen et al. (Chen et al. 2022) employ RL to learn rule mining within KGs. These approaches highlight the benefits of integrating KGs with RL but do not focus on the human-like memory systems or the partially observable environments that our work addresses.
6 Conclusion
In this paper, we introduced an approach to solving POMDPs by leveraging KG-based human-like memory systems. We developed the Rooms Environment with varying levels of complexity, where agents navigate and answer questions based on partial observations. We compared a baseline agent with a simple memory system to our agent, which incorporates episodic and semantic memory systems. Results show that the HumemAI agent outperforms the baseline, effectively capturing information about the belief state and improving decision-making capabilities.
There is room for improvement in our paper. Potential areas for enhancement include jointly learning memory management and exploration policies, which could yield better performance. Additionally, exploring other function approximators, such as graph neural networks (GNNs), might offer improvements over the currently used LSTM. Transitioning from heuristic-based question-answering functions (the reward function) to learning these functions directly could also be beneficial. Furthermore, applying this method to other POMDP environments could test its generalizability and robustness across different scenarios.
7 Acknowledgments
This research was (partially) funded by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organization for Scientific Research, https://www.hybrid-intelligence-centre.nl/.